OpenStudy (anonymous):

FAN & MEDAL! Plz help 3.    Which translation rule describes the translation that is 6 units to the left and 5 units up?  (1 point) (x, y)   (x + 6, y – 5) (x, y)   (x – 6, y – 5) (x, y)   (x + 6, y + 5) (x, y)   (x – 6, y + 5) 4.    Which rule describes the transformation that is a reflection across the y­axis?  (1 point) (x, y) → (x, –y) (x, y) → (–x, y) (x, y) → (–x, –y) (x, y) → (y, x)

OpenStudy (amorfide):

First of all, let's look at the line y=x if you were to translate it in the x axis you would replace x with (x+a) if you were to translate in the y axis, you would simply add a constant C to the equation for example translating y=x 3 units in the negative x and 2 units in the positive y would be y=(x+3)+2

OpenStudy (anonymous):

(x, y) (x - 6, y + 5) y=-x ||| y=x (x, y) → (–x, y)

OpenStudy (anonymous):

Thanks